# A LINAC-based, compact x-ray source by inverse Compton scattering

### v1

@@ 5.1 Short pulse lengths
By changing the arrival time of the laser relative to the arrival time of the electron bunches, the x-ray pulse length can be estimated. Both electron and laser beams were focused down to $\sigma_{e}=\sigma_{L}\approx 20$ µm, and the laser pulse energy was set to $5.0$ mJ. The Rayleigh length of the laser was $z_{R}=6.3$ mm. Figure 6 shows the measured $N_{X}$ as function of delay time with a Gaussian fit, yielding a rms overlay time $\tau_{\mathrm{overlay}}=2.8$ ps.
Because the laser pulse length is significantly smaller than the electron pulse length ($\tau_{L}\ll\tau_{e}$), $\tau_{\mathrm{overlap}}$ is the result of the convolution between the experienced Gaussian laser profile and electron profile. These measurements were performed with $\theta_{L}=(6\pm 2)^{\circ}$, as visualized in the inset in figure 6. The electrons thus experience a Gaussian laser profile with standard deviation $\tau_{\mathrm{int}}=\sigma_{\mathrm{int}}/c=\sigma_{L}/(c\sin\theta_{L})=(0.7\pm 0.3)$ ps. The x-ray pulse length $\tau_{X}$ is equal to $\tau_{e}$, which is found by $\tau_{X}=\tau_{e}=\sqrt{\tau_{\mathrm{overlap}}^{2}-\tau_{\mathrm{int}}^{2}}=(2.3\pm 0.3)$ ps. The uncertainty of the fit was negligible compared to the uncertainty in $\theta_{L}$. An exact value of $\tau_{e}$ strongly depends on $\theta_{L}$ which is experimentally difficult to measure precisely. The results from figure 6 do however confirm a $\tau_{X}$ in the order of picoseconds, with $2.6$ ps as upper bound.
@@ 5.2 Continuous energy tunability
To determine the spectral angular density distribution and energy spectrum of the x-ray beam we have used a hyperspectral x-ray camera [47]. An example of the measured spectral angular density distribution corresponding to 10 MW average input power supplied to the accelerator is shown in figure 7. As reference, figure 7 shows the expected x-ray energy per radial angle based on equation (1). For the reference, $\beta$ was found using the x-ray energy corresponding to the peak of the energy spectrum. The behavior of the measured distribution and equation (1) for different angles $\theta_{X}$ are comparable. Figure 7 furthermore shows that the x-ray energy spread is uncorrelated with respect to $\theta_{X}$ for $\theta_{x}<1.5$ mrad. The following results are therefore based on data acquired within an opening angle $\theta_{x}=1.5$ mrad.
Figure 7: The measured spectral angular density distribution for acceleration with 10 MW average power RF pulses supplied to the LINAC. The white line is a reference for the expected shape based on equation (1).
Figure 8a shows the normalized x-ray energy spectra that has been measured for different average RF input powers supplied to the accelerator, varying the electron energy and thus the x-ray energy. The relation between the average RF input power and the x-ray energy makes it possible to tune the energy in a continuous manner. The spectra before normalization can be found in Supplement 1 (Section S1). The resolution of the used hyperspectral x-ray camera is approximately 152 eV (at the manganese Kα line) [46].
To verify the trend of the generated x-ray spectra as a function of the RF power, the peak x-ray energies of the spectra were plotted for the corresponding average input powers. A fit of the expected trend based on equation (1) is shown alongside the measured data presented in figure 8b. The fit parameter corresponded to $a=6.2\cdot 10^{-2}$ MW. For the lower peak x-ray energies, the average RF input power is higher than expected. This could be attributed to the fact that acceleration at lower average RF input powers is less efficient. Overall, the trend between x-ray energy and RF power is as expected.
Lastly, we have determined the FWHM of the normalized energy spectra presented in figure 8a. The results are shown in figure 8c, illustrating the monochromaticity of the x-ray beam. The FWHM of the different spectra are in the range of 250 eV - 400 eV, which corresponds to 3.6% - 4.5% when normalized to the respective peak energies. The discrepancy of the data point at 8.5 MW compared to the rest could be attributed to instabilities during the measurement. The effect of the opening angle $\theta_{X}$ on the FWHM for 10 MW average RF input power can be found in Supplement 1 (Section S1).
@@ 5.3 Polarization
The polarization of the x-ray photons is directly related to the polarization of the incident laser, which allows for fast polarization control [37, 27]. To verify this ability, the spatial x-ray distribution was measured for different angles of the incident laser polarization $\nu$. In the case of linear laser polarization, according to equation (3) we expect an elliptical-shaped distribution, which rotates with the changing polarization.
We have determined the orientation of the major axis and its angle $\delta$ with the horizontal axis of the detector using Principle Component Analysis (PCA). The orientation angle is plotted for the different polarization angles $\nu$ of the incident laser in figure 9. The figure shows the linear relation between the two variables, which is as expected. Because it was experimentally difficult to manually specify the orientation of the half-wave plate, the 95% prediction interval was added to the plot. The insets show examples of the spatial distribution with lines to guide the eye. The results clearly show a rotation in the spatial distribution, confirming a characteristic feature of ICS. Although the polarization of the x-rays is not directly measured, the spatial distributions indirectly indicate the ability to control the polarization. This control is not restricted to linear polarization, as in the measurements, but can also easily be extended to circular polarization.
@@ 6 Conclusion and outlook
Although the setup can already be used for specific experiments, its development is far from complete. By conditioning the HPRF system the repetition rate can be increased from 50 Hz to 500 Hz. Additionally, the system can be conditioned to produce 200 ns, 24 MW pulses, corresponding to acceleration gradient of 60 MV/m. This will allow for the generation of x-ray energies up to 20 keV. Furthermore, with the current setup the bunch charge and laser pulse energy can be increased to 10 pC and 12 mJ respectively, which will improve the x-ray flux. The flux can be further improved by implementing quadrupoles and an optical telescope, reducing the electron and laser spot sizes. By extrapolating the first results with the future upgrades, the expected average brilliance according to equation (4) is estimated at $3\cdot 10^{8}$ photons/(s $\times$ mrad2 $\times$ mm2 $\times$ 0.1% BW).
The modularity of the Smart*Light setup however allows for a different electron source. Implementing a thermionic gun operating in burst mode would further increase the expected average brilliance to $10^{12}$ photons/(s $\times$ mrad2 $\times$ mm2 $\times$ 0.1% BW) [34]. If high coherence is desired over flux, an ultracold electron source would be suitable [35]. Furthermore, by applying second harmonic generation, the wavelength of the interaction laser can be halved, resulting in a maximum x-ray photon energy of 40 keV.

### v2

@@ 5.1 Short pulse lengths
By changing the arrival time of the laser relative to the arrival time of the electron bunches, the x-ray pulse length can be estimated. Both electron and laser beams were focused down to $\sigma_{e}=\sigma_{L}\approx 20$ µm, and the laser pulse energy was set to $5.0$ mJ. The Rayleigh length of the laser was $z_{R}=6.3$ mm. Fig. 6 shows the measured $N_{X}$ as function of delay time with a Gaussian fit, yielding a rms overlay time $\tau_{\mathrm{overlay}}=2.8$ ps.
Because the laser pulse length is significantly smaller than the electron pulse length ($\tau_{L}\ll\tau_{e}$), $\tau_{\mathrm{overlap}}$ is the result of the convolution between the experienced Gaussian laser profile and electron profile. These measurements were performed with $\theta_{L}=(6\pm 2)^{\circ}$, as visualized in the inset in Fig. 6. The electrons thus experience a Gaussian laser profile with standard deviation $\tau_{\mathrm{int}}=\sigma_{\mathrm{int}}/c=\sigma_{L}/(c\sin\theta_{L})=(0.7\pm 0.3)$ ps. The x-ray pulse length $\tau_{X}$ is equal to $\tau_{e}$, which is found by $\tau_{X}=\tau_{e}=\sqrt{\tau_{\mathrm{overlap}}^{2}-\tau_{\mathrm{int}}^{2}}=(2.3\pm 0.3)$ ps. The uncertainty of the fit was negligible compared to the uncertainty in $\theta_{L}$. An exact value of $\tau_{e}$ strongly depends on $\theta_{L}$ which is experimentally difficult to measure precisely. The results from Fig. 6 do however confirm a $\tau_{X}$ in the order of picoseconds, with $2.6$ ps as upper bound.
@@ 5.2 Continuous energy tunability
To determine the spectral angular density distribution and energy spectrum of the x-ray beam we have used a hyperspectral x-ray camera [48]. An example of the measured spectral angular density distribution corresponding to 10 MW average input power supplied to the accelerator is shown in Fig. 7. As reference, Fig. 7 shows the expected x-ray energy per radial angle based on Eq. (1) (white solid curve). For the reference, $\beta$ was found using the x-ray energy corresponding to the peak of the energy spectrum. The behavior of the measured distribution and Eq. (1) for different angles $\theta_{X}$ are comparable. Fig. 7 furthermore shows that the x-ray energy spread is uncorrelated with respect to $\theta_{X}$ for $\theta_{x}<1.5$ mrad. The following results are therefore based on data acquired within an opening angle $\theta_{x}=1.5$ mrad.
Figure 7: The measured spectral angular density distribution for acceleration with 10 MW average power RF pulses supplied to the LINAC. The white line is a reference for the expected shape based on Eq. (1).
Fig. 8a shows the normalized x-ray energy spectra that has been measured for different average RF input powers supplied to the accelerator, varying the electron energy and thus the x-ray energy. The relation between the average RF input power and the x-ray energy makes it possible to tune the energy in a continuous manner. The spectra before normalization can be found in Supplement 1 (Section S1). The resolution of the used hyperspectral x-ray camera is approximately 144 eV (FWHM at 6 keV) [50].
To verify the trend of the generated x-ray spectra as a function of the RF power, the peak x-ray energies of the spectra were plotted for the corresponding average input powers. A fit of the expected trend based on Eq. (1) is shown alongside the measured data presented in Fig. 8b. The fit parameter corresponded to $a=6.2\cdot 10^{-2}$ MW. For the lower peak x-ray energies, the average RF input power is higher than expected. This could be attributed to the fact that acceleration at lower average RF input powers is less efficient. Overall, the trend between x-ray energy and RF power is as expected.
Lastly, we have determined the FWHM of the normalized energy spectra presented in Fig. 8a. The results are shown in Fig. 8c, illustrating the monochromaticity of the x-ray beam. The FWHM of the different spectra are in the range of 250 eV - 400 eV, which corresponds to 3.6% - 4.5% when normalized to the respective peak energies. The discrepancy of the data point at 8.5 MW compared to the rest could be attributed to instabilities during the measurement. The effect of the opening angle $\theta_{X}$ on the FWHM for 10 MW average RF input power can be found in Supplement 1 (Section S1).
@@ 5.3 Polarization
The polarization of the x-ray photons is directly related to the polarization of the incident laser, which allows for fast polarization control [37, 27]. To verify this ability, the spatial x-ray distribution was measured for different angles of the incident laser polarization $\nu$. In the case of linear laser polarization, according to Eq. (3) we expect an elliptical-shaped distribution, which rotates with the changing polarization.
We have determined the orientation of the major axis and its angle $\delta$ with the horizontal axis of the detector using Principle Component Analysis (PCA). The orientation angle is plotted for the different polarization angles $\nu$ of the incident laser in Fig. 9. The Fig. shows the linear relation between the two variables, which is as expected. Because it was experimentally difficult to manually specify the orientation of the half-wave plate, the 95% prediction interval was added to the plot. The insets show examples of the spatial distribution with lines to guide the eye. The results clearly show a rotation in the spatial distribution, confirming a characteristic feature of ICS. Although the polarization of the x-rays is not directly measured, the spatial distributions indirectly indicate the ability to control the polarization. This control is not restricted to linear polarization, as in the measurements, but can also easily be extended to circular polarization.
@@ 6 Conclusion and outlook
Although the setup can already be used for specific experiments, its development is far from complete. By conditioning the HPRF system the repetition rate can be increased from 50 Hz to 500 Hz. Additionally, the system can be conditioned to produce 200 ns, 24 MW pulses, corresponding to an acceleration gradient of 60 MV/m. This will allow for the generation of x-ray energies up to 20 keV. Furthermore, with the current setup the bunch charge and laser pulse energy can be increased to 10 pC and 12 mJ respectively, which will improve the x-ray flux. The flux can be further improved by implementing quadrupoles and an optical telescope, reducing the electron and laser spot sizes. By extrapolating the first results with the future upgrades, the expected average brilliance according to Eq. (4) is estimated at $3\cdot 10^{8}$ photons/(s $\times$ mrad2 $\times$ mm2 $\times$ 0.1% BW).
The modularity of the Smart*Light setup allows for a different electron source. Implementing a thermionic gun operating in burst mode would further increase the expected average brilliance to $10^{12}$ photons/(s $\times$ mrad2 $\times$ mm2 $\times$ 0.1% BW) [34]. If high coherence is desired over flux, an ultracold electron source would be suitable [35]. Furthermore, by applying second harmonic generation, the wavelength of the interaction laser can be halved, resulting in a maximum x-ray photon energy of 40 keV.